OpenStudy (chupacabraj):

Did I do this right? Finding the surface area of curve y= x^2 (0,1) rotated about the y-axis.

OpenStudy (chupacabraj):

\[SA= 2\pi \int\limits_{0}^{1}x(2x+1) dx\]

OpenStudy (chupacabraj):

\[SA= 2\pi \int\limits_{0}^{1}2x^2 + x dx\]

OpenStudy (chupacabraj):

\[SA= \frac{ 14\pi }{ 6 }\]

OpenStudy (chupacabraj):

I simplified the arc length part

jimthompson5910 (jim_thompson5910):

I'm not sure how you got the 2x+1 part. Can you show your work on that? Hopefully you found that y = x^2 dy/dx = 2x [dy/dx]^2 = 4x^2 [dy/dx]^2+1 = 4x^2+1 agreed?

OpenStudy (chupacabraj):

yes I agreed. I believe I made the mistake of assuming you could easily take the square root of that.

OpenStudy (chupacabraj):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{4x^2+1} \ne \sqrt{4x^2} + \sqrt{1}\]

OpenStudy (chupacabraj):

Ohh ok so plugging into the SA equation I have\[SA= 2\pi \int\limits_{0}^{1}x \sqrt{4x^2+1}dx\]

jimthompson5910 (jim_thompson5910):

you have the correct integral now

OpenStudy (chupacabraj):

ohh you can only do that if there weren't added right? like if they were seperate or a multiplication? what was inside the radical

jimthompson5910 (jim_thompson5910):

yes \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

OpenStudy (chupacabraj):

well, anyways moving on, I integrated using U substitution. I ended up with 5pi/4

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

http://www.wolframalpha.com/input/?i=2pi*int(x*sqrt(4x%5E2%2B1),x%3D0..1) 5pi/4 is not the correct answer

OpenStudy (chupacabraj):

How do you integrate that then?

jimthompson5910 (jim_thompson5910):

u = 4x^2+1 du = 8x*dx x*dx = du/8 agreed so far?

OpenStudy (chupacabraj):

yes, that's what I did

jimthompson5910 (jim_thompson5910):

\[\Large 2\pi\int x\sqrt{4x^2+1}dx\] \[\Large 2\pi\int (\sqrt{4x^2+1})*(x*dx)\] \[\Large 2\pi\int \sqrt{u}*\frac{du}{8}\] \[\Large 2\pi*\frac{1}{8}\int u^{1/2}du\] \[\Large \frac{\pi}{4}\int u^{1/2}du\] what comes next?

OpenStudy (chupacabraj):

ohh I might of forgotten about the 3/2 power at the end :s

OpenStudy (chupacabraj):

pi/4 * 2/3 * u^3/2

jimthompson5910 (jim_thompson5910):

replace the u with 4x^2+1 to get \[\Large \frac{\pi}{4}*\frac{2}{3}*(4x^2+1)^{3/2}+C\]

OpenStudy (chupacabraj):

I get 5.85 now

jimthompson5910 (jim_thompson5910):

close but not quite there

jimthompson5910 (jim_thompson5910):

I think you forgot to plug in x = 0 and subtract

OpenStudy (chupacabraj):

5.18

jimthompson5910 (jim_thompson5910):

If \[\Large g(x) = \Large \frac{\pi}{4}*\frac{2}{3}*(4x^2+1)^{3/2}+C\] then what is g(1) equal to?

OpenStudy (chupacabraj):

ohh algebra mistake, I get 5.33 now

jimthompson5910 (jim_thompson5910):

5.33041350026898 which rounds to 5.33 so it looks good

OpenStudy (chupacabraj):

Thanks :)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (chupacabraj):

Now, if I want to find the SA of the same curve rotated about the x axis, the only thing that changes is the radius right?

OpenStudy (chupacabraj):

\[SA= 2\pi \int\limits_{0}^{1}x^2\sqrt{4x^2+1}dx\]

jimthompson5910 (jim_thompson5910):

You'll now use the formula \[\Large SA = 2\pi\int_{a}^{b}y\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]

jimthompson5910 (jim_thompson5910):

yes correct, evaluate what you just wrote

OpenStudy (chupacabraj):

How would I integrate this? by parts?

jimthompson5910 (jim_thompson5910):

Integration by parts u = sqrt(4x^2+1) du = 4x/sqrt(4x^2+1) dv = x^2 v = (1/3)x^3 \[\Large \int u dv = u*v - \int v du\] \[\Large \int x^2\sqrt{4x^2+1} = \sqrt{4x^2+1}*\frac{1}{3}x^3 - \int \frac{1}{3}x^3 \frac{4x}{\sqrt{4x^2+1}}dx\] I'm getting stuck though

jimthompson5910 (jim_thompson5910):

This person is using hyperbolic trig, which your teacher may have not gone over yet http://mymathforum.com/calculus/48434-integral-x-2-sqrt-1-4x-2-a.html

jimthompson5910 (jim_thompson5910):

so it's probably best to use a calculator to get the approx result if you have a TI83, TI84, etc, you can use the function `FnInt` (which stands for Function Integration) http://tibasicdev.wikidot.com/fnint

jimthompson5910 (jim_thompson5910):

or you can use something like wolfram http://www.wolframalpha.com/input/?i=2*pi*int(x%5E2*sqrt(x%5E2%2B1),x%3D0..1)